Which products should be produced? P 90 baht/unit 100 baht/unit 50 unitslwk 100 units/wk Purchase Part 5 baht/unit 5 min/unit 15 min/unit 10 min/unit 5 min/unit 15 min/unit 15 min/unit 15 min/unit 10 minlunit RM 1 20 baht/unit RM 2 20 baht unit RM 3 20 baht/ unit 8 hr 5 day work week. $6,000 week total operating expense

Calculate the maximum profit that can be earned by producing the 2 goods. It can be Fully Q or Fully P or a combination of both.
Answer #1
  • Fully P
  • Fully Q
  • Combination of both

Fully P - Production is 100 units per week; sold at 90 baht per unit

Revenues = 90* 100 = 9000

total operating expenses = $6000 per week

Profit = $9000 - $6000 = $3000

Q

Number of units produced = 50 units per week

Price = 100 baht per unit

Revenue = $100*50 = 5000

Operating expenses = $6000

The firm makes losses manufacturing only Q.